Lightfoot Inc., a software development firm, has stock outstanding as follows: 25,000 shares of cumulative preferred 2% stock, $20 par, and 31,000 shares of $125 par common. During its first four years of operations, the following amounts were distributed as dividends: first year, $3,750; second year, $6,250; third year, $42,940; fourth year, $88,120. Calculate the dividends per share on each class of stock for each of the four years. Round all answers to two decimal places. If no dividends are paid in a given year, enter "0".

Preferred stock (dividend per share)

1st year:

2nd year:

3rd year:

4th year:

Common stock (dividend per share)

1st year:

2nd year:

3rd year:

4th year:
Answer #1

Calculate dividend per share

Preferred dividend = 25000*20*2% = 10000

Dividend distribution

Total dividend Preferred dividend Common Dividend
1st year 3750 0.15 0
2nd year 6250 0.25 0
3rd year 42940 20000/25000 = 0.80 22940/31000 = 0.74
4th year 88120 10000/25000 = 0.40 78120/31000 = 2.52